A company issues a​ ten-year bond at par with a coupon rate of 6.6​% paid​ semi-annually. The YTM at the beginning of the third year of the bond​ (8 years left to​ maturity) is 9.2​%. What is the new price of the​ bond?

A. $ 855

B. $ 1,026

C. $ 1,197

D. ​$1,000
Answer #1

Price of a bond is mathematically represented as: 1 1-- (1 + i) PEC. - +- м. (1 + i) where P is price of bond, with periodi

M = $1000, n = 8 * 2 = 16 semi-annual periods, i = 9.2%/2 = 4.6% (semi-annually), C = 6.6% *$ 1000/2 = $33 (semi-annually)

1- P = 33* - (1+0.046)16 0.046 + 1000 (1 + 0.046) 16

P = 368.05 + 486.96

P = $855.01(Option A)